# The Coverage Principle: A Framework for Understanding Compositional Generalization

### v1

@@ 4 Experimental setup
Figure 2: Four synthetic compositional structures we study.
Dataset construction
@@
We construct four synthetic compositional tasks: 2-Hop, Parallel 2-Hop, 3-Hop, and Non-tree (Fig. 2). We create random mappings from the product space of a token set to control generalization sources not attributable to compositional structures (i.e., commutativity). [^6] Such external generalization sources will be discussed in Sec. 7 and App. H. We explain the dataset construction process using 2-Hop task (Fig. 2(a)), $(x_{1},x_{2},x_{3})\mapsto t$ with $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, as an example. We construct training datasets by defining a token set with size $|{\mathcal{X}}|$, and creating two random maps for the primitive functions $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$. We mark a fraction $p_{\mathrm{seen}}=0.7$ of each function’s domain as ‘seen’, gather all possible combinations where both functions are applied to inputs from their seen domains, and uniformly sample $N$ examples to form a training dataset. See Sec. B.1 for more details of the dataset construction process.
Training & evaluation
@@
Following Wang et al. [19], we train randomly initialized GPT-2 [75] models with 8 layers, 12 heads, and 768 dimensions (see Sec. B.2 for details). We construct two evaluation sets, each with 2,000 instances: (1) In-Domain (ID) Test Set : all primitive function applications (e.g., $f_{1}(x_{1},x_{2})$ and $f_{2}(b,x_{2})$ in 2-Hop task) are observed during training, but the specific combination was unseen. (2) Out-of-Domain (OOD) Test Set: at least one primitive function application is never observed during training.
5 Results
5.1 Predictive power of $k$-coverage on Transformer’s compositional generalization
@@
Figure 3: Left: Percentage of covered ID data depending on $k$ values and dataset size ($N$), for 2-Hop task ($|{\mathcal{X}}|=50$). Right: Test accuracy depending on $k$-cutoff values for 2-Hop task ($|{\mathcal{X}}|$=50, $N$=10k). Each line represents a different training checkpoint. Note that OOD accuracy remains at chance level (=1/50) regardless of training time. The bars below show the number of test data for each $k$-cutoff value.
@@
We now empirically validate the predictive power of the coverage principle by quantifying the correlation between $k$-coverage and model generalization performance. To this end, we implement and release a task-agnostic coverage determination algorithm (see App. C for implementation details) that can be applied to diverse compositional structures. Then, we analyze what fraction of ID test data of 2-Hop task with $|{\mathcal{X}}|=50$ lies inside $k$-coverage, depending on $k$ and dataset size $N$. Fig. 3 (Left) shows that at $N=5$k, every ID test example is already covered with minimal evidence ($k=1$).
@@
Ideally, a single witness of functional equivalence could suffice. However, experiments show that such minimal coverage alone is insufficient for efficient generalization in practice. To demonstrate this, we define a sample’s $k$-cutoff as the lowest $k$ for which an input lies in $k$-coverage, measuring the strength of evidence for functional equivalence. For example, a $k$-cutoff of 3 means that an example is inside coverage with $k=3$ but not with $k=4$. For 2-Hop dataset with $N=10$k, we classify each ID test instance according to its $k$-cutoff, and track accuracy development for each group across 50k training epochs. As shown in Fig. 3 (Right), generalization success shows a strong relationship with $k$-cutoff values. Test data with low $k$-cutoff values show delayed improvement even after extensive training, while examples with stronger evidence generalize much faster. Notably, OOD test examples (outside coverage) remain at chance-level ($1/50=0.02$) throughout training, empirically validating our coverage principle.
@@
These results yield two important insights. First, successful generalization in practice requires robust coverage so the model can confidently identify and utilize functional equivalence relationships. The $k$ parameter effectively quantifies this evidence strength, directly impacting generalization speed and reliability. Second, while our experiments use uniformly distributed data, the coverage principle explains why models struggle with generalizing long-tail distribution in imbalanced real-world data [76, 77, 78]. Rare combinations naturally receive limited functional equivalence evidence (low $k$), placing them effectively outside practical coverage, despite technically being in-distribution.
5.2 The clustering of latent representation drives generalization on coverage
@@
Next, we investigate how the model internally represents functional equivalence for $k$-covered inputs. Specifically, we inspect a GPT-2 trained on 2-Hop task ($|{\mathcal{X}}|=50$, $N=10\text{k}$) for 50k epochs (corresponding to the yellow line in Fig. 3 (Right)). [^7] Analyses for varying factors including task structures, entity set size ($|{\mathcal{X}}|$), dataset size ($N$), and training steps give consistent results; see App. D. We observe that when a model successfully generalizes to ID test data, it maps functionally equivalent components into tight latent clusters, thereby encoding the equivalence relationships needed for compositional generalization.
@@
To quantify this representation clustering phenomenon, we develop a metric that captures how distinctly the model separates functionally equivalent fragments from others. Specifically, we measure the difference between the average pairwise cosine similarity of latent vectors that share the same intermediate state $b=f_{1}(x_{1},x_{2})$ ($\overline{\cos}_{\text{intra}}$), and those that do not ($\overline{\cos}_{\text{inter}}$), for each position and layer of the model. We term this difference the Intra–Inter Cosine Gap $\text{IICG}=\overline{\cos}_{\text{intra}}-\overline{\cos}_{\text{inter}}$, where higher values indicate stronger within-group clustering relative to between-group separation. Fig. 4 (Left) reveals a clear relationship: higher $k$-cutoff values yield higher IICG scores at certain positions, indicating that stronger functional equivalence evidence leads to more coherent internal representations. Critically, OOD examples show no clustering pattern at all, since they lack any functional equivalence evidence in the training data, as predicted by the coverage principle. The PCA visualization at position $x_{2}$ and layer 3 (Right) shows this trend visually. We verify that the representation clusters play a causal role in ID generalization with causal tracing [67, 66], a widely used mechanistic interpretability technique to identify Transformer circuits (see Fig. 8 in App. D).
@@
Our findings extend the insights of Wang et al. [19] in several ways. First, we demonstrate that unified circuit formation is driven by functional equivalence evidence in the training data, not by explicit exposure to intermediate computation steps. Consequently, out-of-coverage inputs lack these unified circuits by construction, explaining their systematic failure on OOD examples. Moreover, we find that these clustered representations are not necessarily aligned with vocabulary embeddings, implying that standard interpretability methods like logit lens [70] may fail to detect these functional equivalence representations despite their presence. See App. I for a detailed analysis.
5.3 When does coverage lead to full ID generalization? A power-law lower bound
Figure 5: Left: Log-log plot of measured $\hat{N}_{\mathrm{req}}$ vs. token set size ($|{\mathcal{X}}|$) across three compositional tasks. The slope $c$ corresponds to the empirical power-law scaling exponent. Omitted points for 3-Hop are due to prohibitively large dataset requirements. Right: Power-law scaling remains invariant across GPT-2 model sizes (68M to 1.5B parameters) for 2-Hop task. $R^{2}>0.99$ for all linear fitting.
@@ 5.3 When does coverage lead to full ID generalization? A power-law lower bound
Our analysis of $k$-coverage and representation clustering demonstrates that stronger functional equivalence evidence leads to better generalization. A natural follow-up question arises: How large should the training set be to enable full generalization on all ID test data? Intuitively, this requires the training set to support the functional equivalence of every pair of inputs that shares the same intermediate state $b$. Formally, for a 2-Hop task we need $(x_{1},x_{2})\equiv^{\{1,2\}}_{D}(x^{\prime}_{1},x^{\prime}_{2})$ whenever $f_{1}(x_{1},x_{2})=f_{1}(x^{\prime}_{1},x^{\prime}_{2})$. Assuming a learner requires at least $k$ distinct pairs of evidence to establish functional equivalence of two fragments, how does the required dataset size scale with the token set size $|{\mathcal{X}}|$? In practical terms, this question addresses how much data is fundamentally needed for compositional generalization, which is central to understanding data scaling requirements for compositional tasks. For 2-Hop task, we derive the following result (full derivation in App. E):
Result 5.1 (Power-law lower bound).
Let $f:{\mathcal{X}}^{3}\!\to\!{\mathcal{X}}$ be a 2-Hop composition. Assume a learner recognizes functional equivalence of two subsequences only after observing them in at least $k$ distinct pairs of evidence (i.e., functionally $k$-equivalent). [^8] Note that this assumption is equivalent to modeling Fig. 3 (Right) as a step function. Let $N_{\text{req}}(|{\mathcal{X}}|,k)$ be the smallest training dataset size that enables complete generalization to ID examples under this evidence threshold. Then, up to poly-logarithmic factors in $|{\mathcal{X}}|$, $N_{\text{req}}(|{\mathcal{X}}|,k)=\tilde{\Omega}\bigl{(}|{\mathcal{X}}|^{\,\alpha(k)}\bigr{)},\text{where}\;\alpha(k)=2.5-\frac{0.5}{k}.$
@@
Res. 5.1 predicts that a learner relying on pattern matching requires training dataset size scaling at least quadratically with respect to the token set size, to fully generalize on ID test data. To empirically confirm this, we define a practical threshold $\hat{N}_{\mathrm{req}}$ to estimate $N_{\mathrm{req}}(|{\mathcal{X}}|,k)$, as a minimal amount of training data required to exceed ID accuracy of 0.99 within 100 epochs after reaching the same level on training data. [^9] See App. F measurement details. Fig. 5 (Left) shows the measured power-law exponents for $\hat{N}_{\text{req}}$ vs. $|{\mathcal{X}}|$ across different task structures. The measured exponent for 2-Hop ($c=2.26$) aligns well with our theoretical predictions of at least quadratic scaling. Although we derive the theoretical bound only for 2-Hop, we observe clear power-law relationships for more complex structures as well. The higher exponents for Parallel-2-Hop ($c=2.43$) and 3-Hop ($c=2.58$) tasks suggest that extra computational steps essentially add another dimension of relationships that require robust coverage, driving the steeper power-law scaling.
Strikingly, these exponents remain invariant across three different GPT-2 model sizes spanning a 20x range in parameters (from 68M to 1.5B) for all three tasks (Fig. 5 (Right) and Tab. 2 in App. F). This supports that the scaling relationship is primarily determined by data properties rather than model capacity, reinforcing our framework’s data-centric perspective. The observed scaling relationships are robust across different hyperparameters (weight decay and learning rate) and empirical decision criteria for $\hat{N}_{\mathrm{req}}$ (see App. F). Overall, these results suggest that compositional generalization in Transformers requires dataset sizes scaling as a power-law with token set size. This has significant implications for how we approach complex reasoning tasks, suggesting that data curation strategies informed by coverage considerations may be more effective than raw parameter scaling.
5.4 Path ambiguity hinders generalization and interpretability in Transformers
Identical Intermediate StatesFunctional Equivalence$(x_{1},x_{2})$$(x_{1}^{\prime},x_{2})$$(x_{1},x_{2}^{\prime})$$(x_{1}^{\prime},x_{2}^{\prime})$$(x_{1},x_{2}^{\prime\prime})$$(x_{1}^{\prime},x_{2}^{\prime\prime})$
Figure 6: Left: In Non-tree task, the coverage principle predicts that representations of input subsequences with the same intermediate state $b=f_{1}(x_{1},x_{2})$ split into multiple context-dependent state representations, conditioned on $x_{2}$ value. Middle: ID test accuracy after standard training (100 epochs post training accuracy$>$0.99), showing Non-tree task significantly underperforms 2-Hop task ($|{\mathcal{X}}|=50$). Right: IICG heatmap from a model that achieved near-perfect ID accuracy (0.96) after extended training (36k epochs, $|{\mathcal{X}}|=50$, $N=50\text{k}$).
@@ 5.4 Path ambiguity hinders generalization and interpretability in Transformers
Many real-world reasoning tasks involve computational structures where a single variable affects the output through multiple paths, creating what we call path ambiguity. Our framework predicts that in such cases, Transformers struggle with forming unified representations of intermediate states that are theoretically equivalent, instead developing context-dependent state representations that vary based on the input context. In this section, we analyze Non-tree task (Fig. 2(d)) as a case study, where $x_{2}$ affects the output through two paths: as input to $f_{1}$ and directly to $f_{2}$. Following the definition of functional equivalence (Def. 3.1), this path ambiguity prevents the model from establishing the functional equivalence of two subsequences $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ that produce the same intermediate state $b$, unless they also share the same $x_{2}$ value ($x_{2}=x_{2}^{\prime}$). Consequently, the model creates context-dependent state representations rather than unifying them to represent true intermediate state $b$, as illustrated in Fig. 6 (Left).
@@
In consequence, the path ambiguity hinders both generalization performance and efficiency, as the model must establish functional equivalence for each $x_{2}$-conditioned equivalent pair. Fig. 6 (Middle) shows that GPT-2 can fully generalize on ID test set of 2-Hop task within a reasonable time with increasing data size, but fails with Non-tree task, even provided with a near-exhaustive amount of possible ID combinations as training data. [^10] For $|\mathcal{X}|=50$ and $p_{\text{seen}}=0.7$, our largest run ($N=50\text{k}$) includes virtually the entire domain ($\approx 0.7^{2}\!\times|\mathcal{X}|^{3}\;\approx\;61\text{k}$ distinct ID triples). Notably, scaling to 1.5B parameters does not show significant improvement in the performance (Fig. 17 in App. G). Extremely prolonged training (36k epochs) with near-exhaustive ID combinations eventually achieves ID accuracy of 0.96. However, IICG analysis reveals no evidence of a unified intermediate state representation formation, with near-zero IICG scores when grouping by the intermediate state value $b$ (Fig. 6 (Right)). In contrast, grouping by $x_{2}$-conditioned intermediate state ($(b,x_{2})$) leads to high IICG scores, suggesting the formation of context-dependent state representations.
This context-dependence raises an interpretability concern, as standard linear probing-based techniques like logit lens [70, 71] would not reliably identify intermediate states. Moreover, our analysis provides insights into why LLMs fall short in planning tasks [29, 31, 30], as planning often requires correct tracking of intermediate states, which can influence outcomes through multiple paths.
5.5 CoT supervision improves data efficiency but still struggles with path ambiguity

### v2

@@ 4 Experimental Setup
Figure 2: Four synthetic task structures we study.
@@
Dataset construction.   We construct four synthetic tasks with different structures: 2-Hop, Parallel 2-Hop, 3-Hop, and Non-tree (Fig.˜2). To isolate functional-equivalence-based generalizations, we create random mappings from a product space of token sets to control the generalization sources not attributable to compositional structures (i.e., commutativity). We explain the dataset construction process using 2-Hop task (Fig.˜2(a)), $(x_{1},x_{2},x_{3})\mapsto t$ with $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, as an example. We construct training datasets by defining a token set with size $\left|{\mathcal{X}}\right|$, and creating two random maps for the primitive functions $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$. We mark a fraction $p_{\mathrm{seen}}=0.7$ of each function’s domain as ‘seen’, gather all possible combinations where both functions are applied to inputs from their seen domains, and uniformly sample $N$ examples to form a training dataset. See Sec.˜B.1 for more details of the dataset construction process.
@@
Training & evaluation.   Following wang2022interpretability, we train randomly initialized GPT-2 (radford2019language) models with 8 layers, 12 heads, and 768 dimensions as a base model (see Sec.˜B.2 for details). We construct two evaluation sets, each with 2,000 instances: (1) ID Test Set : all primitive function applications (e.g., $f_{1}(x_{1},x_{2})$ and $f_{2}(b,x_{3})$ in 2-Hop task) are observed during training, but their specific combination was unseen. (2) Out-of-coverage (canonical OOD) Test Set: at least one primitive function application is never observed during training, which is used as a control group.
5 Quantitative Analysis of Pattern Matching in Transformers
5.1 Evidence Strength is Tightly Aligned to Pattern-Matching Success
@@
Figure 3: Left: Percentage of covered ID data depending on $k$ values and dataset size ($N$), for 2-Hop task ($\left|{\mathcal{X}}\right|=50$). Right: Test accuracy depending on $k$-cutoff values for 2-Hop task ($\left|{\mathcal{X}}\right|=50$, NN=10k). Each line represents a different training checkpoint. Note that out-of-coverage ($k=0$) accuracy remains at chance level ($\approx 1/50$) regardless of training time. The bars below show the number of test data for each $k$-cutoff value.
@@
We first analyze the correlation between $k$-coverage and ID generalization performance of the GPT-2 model. To this end, we implement and release a task-agnostic coverage determination algorithm (see App.˜C) that can be applied to diverse compositional structures. Then, we analyze what fraction of ID test data of 2-Hop task with $\left|{\mathcal{X}}\right|=50$ lies inside $k$-coverage, depending on $k$ and dataset size $N$. Fig.˜3 (Left) shows that at N=5N=5k, every ID test example is already covered with minimal evidence ($k=1$). Hence, in an ideal scenario where a single witness of functional equivalence suffices, training with the dataset as small as N=5N=5k will lead to perfect ID generalization.
@@
However, we demonstrate in our experiments that minimal coverage (i.e., $k=1$) alone is practically insufficient for ID generalization. To demonstrate this, let us fix a training dataset $D$ and define the $k$-cutoff of each input sequence as the lowest value of $k$ for which an input lies in $k$-coverage, measuring the strength of evidence for functional equivalence. For example, a $k$-cutoff of 3 means that an example is inside coverage with $k=3$ but not with $k=4$. For out-of-coverage data, we define $k$-cutoff as 0. Then, for the 2-Hop dataset with N=10N=10k, we classify each ID test instance according to its $k$-cutoff, and track the accuracy development of the GPT-2 model for each group across 50k training epochs. As shown in Fig.˜3 (Right), ID test accuracy shows a strongly positive correlation with $k$-cutoff values. Test data with low $k$-cutoff values show delayed improvement even after extensive training, while examples with stronger evidence generalize much faster.
@@
These results yield two important insights. First, successful ID generalization in practice requires a robust coverage so the model can confidently identify and utilize functional equivalence relationships. The parameter $k$ effectively quantifies this evidence strength, directly impacting generalization speed and reliability. Second, while our experiments use uniformly sampled datasets, the results can explain why models struggle with generalizing long-tail distributions in imbalanced real-world data (mallen2022not; kandpal2023large; chang2024large). Despite technically being in-distribution, rare combinations naturally receive limited evidence of functional equivalence (low $k$), effectively and practically placing them outside the coverage. We believe that our insights will motivate future research on targeted data augmentation strategies to maximize $k$-coverage.
5.2 Latent Representation Clusters Drive Pattern Matching on Coverage
@@
Next, we investigate how the model internally represents functional equivalence for $k$-covered inputs. Specifically, we inspect a GPT-2 trained on 2-Hop task ($\left|{\mathcal{X}}\right|=50$, $N=10\text{k}$) for 50k epochs (corresponding to the yellow line in Fig.˜3 (Right)). [^4] The analyses for varying factors including task structures, entity set size ($\left|{\mathcal{X}}\right|$), dataset size ($N$), and training steps give consistent results; see App. D. We observe that when a model successfully generalizes to ID test data, it maps functionally equivalent components into tight latent clusters, thereby encoding the equivalence relationships needed for compositional generalization.
@@
To quantify this representation clustering phenomenon, we develop a metric that captures how distinctly the model separates functionally equivalent fragments from others. Specifically, we measure the difference between the average pairwise cosine similarity of latent vectors that share the same intermediate state $b=f_{1}(x_{1},x_{2})$ ($\overline{\cos}_{\text{intra}}$), and those that do not ($\overline{\cos}_{\text{inter}}$), for each position and layer of the model. We term this difference the intra–inter cosine gap $\text{IICG}=\overline{\cos}_{\text{intra}}-\overline{\cos}_{\text{inter}}$, where higher values indicate stronger within-group clustering relative to between-group separation. Fig.˜4 (Left) reveals a positive correlation: higher $k$-cutoff values yield higher IICG scores at certain positions, indicating that stronger functional equivalence evidence leads to more coherent internal representations. In contrast, out-of-coverage ($k=0$) examples exhibit no clustering pattern, as they lack evidence of functional equivalence in the training data. The PCA visualization at position $x_{2}$ and layer 3 (Right) shows this trend visually. We verify that the representation clusters play a causal role in pattern matching with causal tracing (goldowsky2023localizing; hanna2023does), a widely used technique to identify Transformer circuits (Fig.˜8).
@@
Our findings extend the previous insights from mechanistic interpretability literature (wang2024grokked) in several ways. First, we demonstrate that unified circuit formation is driven by functional equivalence evidence in the training data, not by explicit exposure to intermediate computation steps. Moreover, we find that these clustered representations are not necessarily aligned with vocabulary embeddings, implying that standard interpretability methods like logit lens nostalgebraist2020logitlens may fail to detect these functional equivalence representations despite their presence (see App.˜E).
6 Data Scaling Law of Pattern Matching Behaviors
@@
Our analysis in the previous section demonstrates that stronger functional equivalence evidence leads to better generalization. A natural follow-up question arises: How large should the training set be to enable full generalization on all ID test data? Intuitively, this requires the training set to support (strongly enough) the functional equivalence of every pair of inputs that shares the same intermediate state $b$. Formally, for a 2-Hop task we need $(x_{1},x_{2})\equiv^{\{1,2\}}_{D}(x^{\prime}_{1},x^{\prime}_{2})$ whenever $f_{1}(x_{1},x_{2})=f_{1}(x^{\prime}_{1},x^{\prime}_{2})$. Assuming that generalization is constrained by the $k$-coverage, how should the train data size scale in the token set size $\left|{\mathcal{X}}\right|$ to achieve it completely? In practical terms, this question seeks to determine the amount of data necessary and/or sufficient to cover all ID combinations with $k$-coverage, which is crucial for understanding the data cost for pattern-matching generalization. To quantify the data cost, we establish and prove the following sample-complexity upper bound for learning a 2-Hop task (the complete statement and proof are provided in App.˜F):
Figure 5: Left: Log-log plot of measured $\hat{N}_{\mathrm{req}}$ vs. token set size ($\left|{\mathcal{X}}\right|$) across three compositional tasks. The slope $c$ corresponds to the empirical power-law scaling exponent. Omitted points for 3-Hop are due to prohibitively large dataset requirements. Right: Power-law scaling behavior on 2-Hop task across varying GPT-2 model sizes (68M to 1.5B parameters) and Mamba model (For Mamba, we used 4 layers, a hidden dimension of 256, and a learning rate of 0.008, and $\hat{N}_{\mathrm{req}}$ is measured for only $\left|{\mathcal{X}}\right|\leq 100$, since a larger token set size led to training instability). $R^{2}>0.99$ for all linear fitting.
Theorem 6.1 (Informal; Corollaries˜F.9 and F.17).
Consider a 2-Hop task with a token set of size $n$. For a uniformly randomly sampled train dataset $D$ of size $N$, consider a learner that generalizes within the $k$-coverage of $D$. Then, for large enough $n$, the learner achieves perfect ID generalization with high probability if $N\!\gtrsim\!n^{c}$ with $c=2.5-\frac{0.5}{k}$. In contrast, the learner (with $k\!\geq\!2$) does not achive perfect ID generalization with high probability for some 2-Hop task if $n^{2}\!\lesssim\!N\!\lesssim\!n^{c}$. Here, we ignore the polylogarithmic factors in $n$.
@@
Theorem 6.1 presents a tight sample complexity bound $\tilde{\Theta}(n^{c})$ for pattern-matching-based generalization, showing that the training dataset required to ensure full ID generalization grows polynomially in the token set size $n$, with an exponent $c\in[2,2.5)$. To empirically confirm this, we define a practical threshold $\hat{N}_{\mathrm{req}}$ to estimate $N_{\mathrm{req}}(\left|{\mathcal{X}}\right|,k)$, as a minimal amount of training data required to exceed ID accuracy of 0.99 within 100 epochs after reaching the same level on training data (see App.˜G for the measurement details). Fig.˜5 (Left) shows the measured power-law exponents for $\hat{N}_{\text{req}}$ vs. $\left|{\mathcal{X}}\right|$ across different task structures. The measured exponent for 2-Hop ($c=2.26$) aligns well with our theoretical predictions. Although we derive the theoretical bound only for 2-Hop, we observe clear power-law relationships for more complex structures as well. The higher exponents for Parallel-2-Hop ($c=2.43$) and 3-Hop ($c=2.58$) tasks suggest that extra computational steps essentially add another dimension of relationships that require robust coverage, driving the steeper power-law scaling.
These exponents remain invariant across three different GPT-2 model sizes spanning a 20x range in parameters (from 68M to 1.5B) for all three tasks (Fig.˜5 Right and Tab.˜2 in App.˜G). We also show that the exponent measured with a Mamba model (4 layers and a hidden dimension of 256) falls inside the boundary predicted by the theory (same figure). Interestingly, the result in Fig.˜6 (Middle) demonstrates that with increasing training dataset size $N$, there is a sharp phase transition from ID generalization failure to complete success near $N=20\text{k}$.
Overall, the results support that the data scaling law is primarily determined by data properties rather than model capacity or architectures, and additional generalization mechanisms will be required to achieve milder scaling laws on such compositional tasks. [^5] The observed scaling relationships are robust across different hyperparameters (weight decay and learning rate) and empirical decision criteria for $\hat{N}_{\mathrm{req}}$ (see App. G). We note that our result aligns with the practical observation that parameter scaling does not significantly improve the multi-hop reasoning capability of LLMs (yang2024large) and the data-hungry nature of compositional tasks (lake2018generalization), suggesting that these could be partly attributed to pattern-matching behaviors. We leave further analysis of the connection between scaling behavior and pattern matching as an exciting future research direction.
7 Path Ambiguity Problem as a Failure Case of Pattern Matching
@@
We identify a path ambiguity problem with our framework, a previously uncharacterized failure mode that pattern matching struggles with task structures where a single variable affects the output through multiple paths. In this section, we analyze Non-tree task (Fig.˜2(d)) as a case study, where $x_{2}$ affects the output through two paths, as input to $f_{1}$ and directly to $f_{2}$. Unlike in the 2-Hop case, one cannot establish the functional equivalence of two subsequences $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ that produce the same intermediate state $b$, unless they also share the same $x_{2}$ value ($x_{2}=x_{2}^{\prime}$). It is because $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ are not guaranteed to behave identically (i.e., $f(x_{1},x_{2},x_{3})$ is not necessarily equal to $f(x_{1}^{\prime},x_{2}^{\prime},x_{3})$) when $x_{1}\neq x_{2}$. Consequently, we can predict that Transformers trained on the Non-tree will create context-dependent state representations that are conditioned on $x_{2}$ values, failing to unify them to represent the true intermediate state $b$ (Fig.˜6 Left).
[…]
Figure 6: Left: In Non-tree task, the representations of input subsequences with the same intermediate state $b=f_{1}(x_{1},x_{2})$ are split into multiple context-dependent state representations, conditioned on $x_{2}$ value. Middle: ID test accuracy after standard training with varying training dataset size ($\left|{\mathcal{X}}\right|=50$, evaluated 100 epochs after training accuracy reaches 0.99). Observe a sharp transition from ID generalization failure to complete success near $N=20\text{k}$ for 2-Hop, which does not occur in Non-tree task. Right: IICG heatmap from a model that achieved near-perfect ID accuracy (0.96) after extended training (36k epochs, $\left|{\mathcal{X}}\right|=50$, $N=50\text{k}$).
@@
Experiments show that the path ambiguity indeed hinders both generalization on the ID test set and the interpretability of intermediate state representations, as the model now establishes functional equivalence for each $x_{2}$-conditioned equivalent pair. Fig.˜6 (Middle) shows that GPT-2 can fully generalize on the ID test set of 2-Hop task within a reasonable time with increasing data size, but fails with Non-tree task, even provided with a near-exhaustive amount of possible ID combinations as training data. [^6] For $\left|{\mathcal{X}}\right|=50$ and $p_{\text{seen}}=0.7$, our largest run ($N=50\text{k}$) includes virtually the entire domain ($\approx 0.7^{2}\!\times\left|{\mathcal{X}}\right|^{3}\;\approx\;61\text{k}$ distinct ID triples). Notably, scaling to 1.5B parameters does not show significant improvement in the performance (Fig.˜18), and the Mamba model used in Sec.˜6 shows the same trend of generalization failure (Fig.˜19). In addition, extremely prolonged training (36k epochs) with near-exhaustive ID combinations eventually achieves ID accuracy of 0.96; however, IICG analysis reveals no evidence of a unified intermediate state representation formation, with near-zero IICG scores when grouping by the intermediate state value $b$ (Fig.˜6 Right). In contrast, grouping by $x_{2}$-conditioned intermediate state $(b,x_{2})$ leads to high IICG scores, showing the formation of context-dependent state representations. This context-dependence due to path ambiguity raises an interpretability concern, as standard linear probing-based techniques like logit lens (nostalgebraist2020logitlens; belrose2023eliciting) would not reliably identify intermediate states when a model relies on pattern matching.
Hence, a generalization mechanism other than pattern matching will be required for a robust ID generalization on complex task structures that require the access and update of intermediate states through multiple paths (e.g., planning tasks (ruis2020benchmark; kambhampati2024position; valmeekam2023planbench)), where further characterization of this problem remains as an exciting future direction.
8 CoT Improves Data Efficiency but Path Ambiguity Persists